Computing Basic and Diluted Earnings per Share

Zeller Corporation began 2015 with 122,000 shares of common stock and 16,000 shares of convertible preferred stock outstanding. On March 1 an additional 10,000 shares of common stock were issued. On August 1, another 16,000 shares of common stock were issued. On November 1, 6,000 shares of common stock were acquired for the treasury. The preferred stock has a $2 per-share dividend rate, and each share may be converted into one share of common stock. Zeller Corporation's 2015 net income is $501,000.

Required
a. Compute basic earnings per share for 2015. Round to two decimal places.
$Answer



b. Compute diluted earnings per share for 2015. Round to two decimal places.
Answer #1

Weighted average shares of Common Stock = (122000*2/12+132000*5/12+148000*3/12+142000*2/12) = 136000

a) Basic earning per share = (net income-preferred dividend)/Share outstanding = (501000-32000)/136000 = 3.45

b) Diluted earning per share = 501000/152000 = 3.30